Whats the y slope form for this using y=mx+bto
1.) x-2y=4
2.)x+2y=3
3.) 3x-y=-5

Answer by checkley79(3341)

1.) x-2y=4
-2Y=-X+4
Y=-X/-2+4/-2
Y=X/2-2 SLOPE=1/2.
2.)x+2y=3
2Y=-X+3
Y=-X/2+3/2 SLOPE=-1/2.
3.) 3x-y=-5
-Y=-3X-5
Y=3X+5 SLOPE=3.